Refer to the table below. Assume that the company involved now finds that the learning rate is 91 percent. The 91 percent learning rate coefficient is −0.136062; that is b = −0.136062 in the formula Y = aXb.

Required:

Using the formula, Y = aXb, compute the labor time and costs for 5, 6, and 7 units. (Round your intermediate calculations (except for coefficient) and final answers to 2 decimal places.)

Unit Produced (X) Labor Time Required to Produce the Xth Unit (i.e., the last single unit produced)1 (Y) Cumulative Total Time in Labor- Hours(2) Total Cost(3) Average Cost per Unit(4)
1 100.00 100.00 $5,000.00 $5,000.00
2 91.00 191.00 9,550.00 4,775.00
3 86.12 277.12 13,855.78 4,618.59
4 82.81 359.93 17,996.28 4,499.07
5
6
7
8 75.36 670.72 33,536.02 4,192.00
Answer #1
Unit Produced (X) Labor Time Required to Produce the Xth Unit (i.e., the last single unit produced)1 (Y) Cumulative Total Time in Labor- Hours(2) Total Cost(3) Average Cost per Unit(4)
1                              100.00                      100.00 $5,000.00 $5,000.00
2                                91.00                      191.00 $9,550.00 $4,775.00
3                                86.12                      277.12 $13,855.78 $4,618.59
4                                82.81                      359.93 $17,996.28 $4,499.07
5                                80.33                      440.26 $22,012.73 $4,402.55
6                                78.37                      518.63 $25,931.18 $4,321.86
7                                76.74                      595.37 $29,768.14 $387.91
8                                75.36                      670.72 $33,536.02 $4,192.00

Detail Working for Easy refrence

Computation of Required detail
At Unit Produced = 5
Labour time required= 100X(5^-0.136062)=80.33
Cummulative Total Time in Labour Hour= 359.93+80.33=440.26
Total Cost= (17996.28/359.93)*440.26= 2201273.09
Average Cost per Unit= 22012.73/5=4402.55
At Unit Produced = 6
Labour time required= 100X(6^-0.136062)=78.37
Cummulative Total Time in Labour Hour= 440.26+78.37=518.63
Total Cost= (22012.73/440.26)*518.63=25931.18
Average Cost per Unit= 25931.18/6=4321.86
At Unit Produced = 7
Labour time required= 100X(7^-0.136062)=76.74
Cummulative Total Time in Labour Hour= 518.63+76.74=595.37
Total Cost= (25931.18/518.63)*595.37=29768.13
Average Cost per Unit= 29768.13/7=4252.59